AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Nov 22 11:32:11 2025
Author
*
:
Subject
*
:
<p>11:45 New run started - run 81, same settings as previous runs (75, 76 etc)</p> <p>ROOT histograms NOT zeroed</p> <p>Beam at 2.5e8, target density at 3e12</p> <p>11:50 All checks okay</p> <p>12:45 All checks okay</p> <p>13:45 All checks okay</p> <p>14:45 All checks okay</p> <p>15:45 All checks okay.</p> <p>17:00 All Checks okay. Target temp manually dropped from 30.4K to 30.1K to try and improve density as would not get much above 2.5-3ish. Also played with target position for a while - now seemed to have stabilised it at 4.5e12 as of now.</p> <p>18:00 All checkd okay. Target been stable at 4.5-5.3e12.</p> <p>18:20 Done checks again (all ok) as the beam is being a little stange. Attatched ss of peaky peaks (some of them very nice) but detector 3 histogram wont cooperate and wont let me see the counts if I zoom further down, but can see proton peak fine. Seeing periodic spikes in the counts plot for the particle detector - monitoring. Also the tail of the beam seems more red than usual.</p> <p>19:00 More photos of strange spikes in the counts for particle detector uploaded. Sometimes, the baseline has decent sized spikes too. Attatchments 10-14</p> <p>19:20 Checks ok. Traget stable. Spikes in particle detector seem to be more frequent, uploading more photos below. Also attatched peaks at this stage (attatchemtn 15).</p> <p>19:36 Very strange spikes in particle det. Two bigish peaks at baseline and then strange beginning at the injection. Attatchment 19.</p> <p>19:45 Another spike in the baseline for particle det. Attatchment 20. </p> <p>20:25 Checks ok. Spikes still in particle det plot.</p> <p>21:20 Checks ok. Screenshot of peaks in attatchment 21.</p> <p>22:20 Checks ok. Target been relativly stable all shift at 5e12. Spikes still periodically turning up in particle detector. Histograms are gaining counts, attatchment 22 is the peaks.</p> <p>22:30 Target suddenly got really dense up to 1e13. Lifetime of the beam is ok. Trending down slowly. Did checks again and all is good.</p> <p> </p> <p>01:45. Checks OK. Target seems thick.</p> <p>04:30</p> <p>Temperatures OK - attach</p> <p>Rates OK - attach</p> <p>Merger OK - attach</p> <p>Hisograms look OK - attached</p> <p>Target looks very thick from ring vacuum worsening.</p> <p>7:15 all checks okay - historgams attached</p> <p>beam 2e8, target 1e13, scintillator rate 2000 in peak</p> <p>10:00 </p> <p>Run stopped to gain cave access.</p> <p>Before beam was stopped, target was moved out. We still see some signal on the neutral particle detector, roughly a factor 10 below with target on at 1e13 at/cm2</p> <p>Will need a beam background run for normalisation procedure. See attachment 29</p> <p>10:30</p> <p>Now we have no beam in the ring we still observe some events in the neutrals detector. Some spikes up to 150 counts observed, could explain some of what we have seen so far. See attachment 30</p> <p>Note we are saving all this data to MBS so it is effectively background for the neutral detector. </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.3-7933898